\begin{thebibliography}{99}

%\cite{Bzdak:2012tp}
\bibitem{Bzdak:2012tp} 
  A.~Bzdak and D.~Teaney,
  %``Longitudinal fluctuations of the fireball density in heavy-ion collisions,''
  Phys.\ Rev.\ C {\bf 87}, no. 2, 024906 (2013)
%  doi:10.1103/PhysRevC.87.024906
  [arXiv:1210.1965 [nucl-th]].
  %%CITATION = doi:10.1103/PhysRevC.87.024906;%%
  %32 citations counted in INSPIRE as of 27 Feb 2017
  
%\cite{Bzdak:2015dja}
\bibitem{Bzdak:2015dja} 
  A.~Bzdak and P.~Bozek,
  %``Multiparticle long-range rapidity correlations from fluctuation of the fireball longitudinal shape,''
  Phys.\ Rev.\ C {\bf 93}, no. 2, 024903 (2016)
%  doi:10.1103/PhysRevC.93.024903
  [arXiv:1509.02967 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevC.93.024903;%%
  %9 citations counted in INSPIRE as of 27 Feb 2017
  
%\cite{Aaboud:2016jnr}
\bibitem{Aaboud:2016jnr} 
  M.~Aaboud {\it et al.} [ATLAS Collaboration],
  %``Measurement of forward-backward multiplicity correlations in lead-lead, proton-lead, and proton-proton collisions with the ATLAS detector,''
  Phys.\ Rev.\ C {\bf 95}, no. 6, 064914 (2017)
%  doi:10.1103/PhysRevC.95.064914
  [arXiv:1606.08170 [hep-ex]].
  %%CITATION = doi:10.1103/PhysRevC.95.064914;%%
  %16 citations counted in INSPIRE as of 10 Sep 2017
  
%\cite{Jia:2015jga}
\bibitem{Jia:2015jga} 
  J.~Jia, S.~Radhakrishnan and M.~Zhou,
  %``Forward-backward multiplicity fluctuation and longitudinal harmonics in high-energy nuclear collisions,''
  Phys.\ Rev.\ C {\bf 93}, no. 4, 044905 (2016)
%  doi:10.1103/PhysRevC.93.044905
  [arXiv:1506.03496 [nucl-th]].
  %%CITATION = doi:10.1103/PhysRevC.93.044905;%%
  %23 citations counted in INSPIRE as of 27 Feb 2017
  
%\cite{Bozek:2015tca}
\bibitem{Bozek:2015tca} 
  P.~Bozek, W.~Broniowski and A.~Olszewski,
  %``Two-particle correlations in pseudorapidity in a hydrodynamic model,''
  Phys.\ Rev.\ C {\bf 92}, no. 5, 054913 (2015)
%  doi:10.1103/PhysRevC.92.054913
  [arXiv:1509.04124 [nucl-th]].
  %%CITATION = doi:10.1103/PhysRevC.92.054913;%%
  %9 citations counted in INSPIRE as of 27 Feb 2017

%\cite{Monnai:2015sca}
\bibitem{Monnai:2015sca} 
  A.~Monnai and B.~Schenke,
  %``Pseudorapidity correlations in heavy ion collisions from viscous fluid dynamics,''
  Phys.\ Lett.\ B {\bf 752}, 317 (2016)
%  doi:10.1016/j.physletb.2015.11.063
  [arXiv:1509.04103 [nucl-th]].
  %%CITATION = doi:10.1016/j.physletb.2015.11.063;%%
  %14 citations counted in INSPIRE as of 27 Feb 2017
  
%\cite{Broniowski:2015oif}
\bibitem{Broniowski:2015oif} 
  W.~Broniowski and P.~Bozek,
  %``Simple model for rapidity fluctuations in the initial state of ultrarelativistic heavy-ion collisions,''
  Phys.\ Rev.\ C {\bf 93}, no. 6, 064910 (2016)
%  doi:10.1103/PhysRevC.93.064910
  [arXiv:1512.01945 [nucl-th]].
  %%CITATION = doi:10.1103/PhysRevC.93.064910;%%
  %6 citations counted in INSPIRE as of 27 Feb 2017
  
%\cite{Schenke:2016ksl}
\bibitem{Schenke:2016ksl} 
  B.~Schenke and S.~Schlichting,
  %``3D glasma initial state for relativistic heavy ion collisions,''
  Phys.\ Rev.\ C {\bf 94}, no. 4, 044907 (2016)
%  doi:10.1103/PhysRevC.94.044907
  [arXiv:1605.07158 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevC.94.044907;%%
  %10 citations counted in INSPIRE as of 27 Feb 2017
  
%\cite{Bzd-Dus}
\bibitem{Bzd-Dus} 
  A.~Bzdak and K.~Dusling,
  %``Probing proton fluctuations with asymmetric rapidity correlations,''
  Phys.\ Rev.\ C {\bf 93}, no. 3, 031901 (2016)
%  doi:10.1103/PhysRevC.93.031901
  [arXiv:1511.03620 [hep-ph]];
  %``Saturation scale fluctuations and multi-particle rapidity correlations,''
  Phys.\ Rev.\ C {\bf 94}, 044918 (2016)
%  doi:10.1103/PhysRevC.94.044918
  [arXiv:1607.03219 [hep-ph]].
  
%\cite{Ke:2016jrd}
\bibitem{Ke:2016jrd} 
  W.~Ke, J.~S.~Moreland, J.~E.~Bernhard and S.~A.~Bass,
  %``Constraints on rapidity-dependent initial conditions from charged particle pseudorapidity densities and two-particle correlations,''
  arXiv:1610.08490 [nucl-th].
  %%CITATION = ARXIV:1610.08490;%%
  %1 citations counted in INSPIRE as of 27 Feb 2017
  
%\cite{He:2017laa}
\bibitem{He:2017laa} 
  R.~He, J.~Qian and L.~Huo,
  %``Analyzing fluctuations in rapidity with two-particle correlations and a toy granular model,''
  arXiv:1702.03137 [nucl-th].
  %%CITATION = ARXIV:1702.03137;%%
  %1 citations counted in INSPIRE as of 27 Feb 2017
  
%\cite{star:a1}
\bibitem{star:a1} 
  S. Jowzaee [STAR Collaboration],
  %``Rapidity correlations in the RHIC Beam Energy Scan,''
  talk at the the XXVI international conference on ultrarelativistic heavy-ion collisions (Quark Matter 2017), Chicago, 2017.
  
%\cite{alice:a1}
\bibitem{alice:a1} 
  Saehanseul Oh  [ALICE Collaboration],
  %``Azimuthal correlations of longitudinal structure at midrapidity in Pb-Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV with ALICE,''
  poster at the the XXVI international conference on ultrarelativistic heavy-ion collisions (Quark Matter 2017), Chicago, 2017.
  
%\cite{Bozek:2010vz}
\bibitem{Bozek:2010vz} 
  P.~Bozek, W.~Broniowski and J.~Moreira,
  %``Torqued fireballs in relativistic heavy-ion collisions,''
  Phys.\ Rev.\ C {\bf 83}, 034911 (2011)
%  doi:10.1103/PhysRevC.83.034911
  [arXiv:1011.3354 [nucl-th]].
  %%CITATION = doi:10.1103/PhysRevC.83.034911;%%
  %59 citations counted in INSPIRE as of 03 Mar 2017
  
%\cite{Bialas:2011xk}
\bibitem{Bialas:2011xk} 
  A.~Bialas and K.~Zalewski,
  %``Multibin long-range correlations,''
  Nucl.\ Phys.\ A {\bf 860}, 56 (2011)
%  doi:10.1016/j.nuclphysa.2011.05.006
  [arXiv:1101.1907 [hep-ph]].
  %%CITATION = doi:10.1016/j.nuclphysa.2011.05.006;%%
  %14 citations counted in INSPIRE as of 27 Feb 2017
  
%\cite{Bialas:2011bz}
\bibitem{Bialas:2011bz} 
  A.~Bialas, A.~Bzdak and K.~Zalewski,
  %``Hidden Asymmetry and Long Range Rapidity Correlations,''
  Phys.\ Lett.\ B {\bf 710}, 332 (2012)
%  doi:10.1016/j.physletb.2012.03.008
  [arXiv:1107.1215 [hep-ph]].
  %%CITATION = doi:10.1016/j.physletb.2012.03.008;%%
  %12 citations counted in INSPIRE as of 27 Feb 2017
  
%\cite{Csernai:2012mh}
\bibitem{Csernai:2012mh} 
  L.~P.~Csernai, G.~Eyyubova and V.~K.~Magas,
  %``New method for measuring longitudinal fuctuations and directed flow in ultrarelativistic heavy ion reactions,''
  Phys.\ Rev.\ C {\bf 86}, 024912 (2012)
  Erratum: [Phys.\ Rev.\ C {\bf 88}, 019902 (2013)]
%  doi:10.1103/PhysRevC.88.019902, 10.1103/PhysRevC.86.024912
  [arXiv:1204.5885 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevC.88.019902, 10.1103/PhysRevC.86.024912;%%
  %19 citations counted in INSPIRE as of 03 Mar 2017
  
%\cite{Pang:2012he}
\bibitem{Pang:2012he} 
  L.~Pang, Q.~Wang and X.~N.~Wang,
  %``Effects of initial flow velocity fluctuation in event-by-event (3+1)D hydrodynamics,''
  Phys.\ Rev.\ C {\bf 86}, 024911 (2012)
%  doi:10.1103/PhysRevC.86.024911
  [arXiv:1205.5019 [nucl-th]].
  %%CITATION = doi:10.1103/PhysRevC.86.024911;%%
  %85 citations counted in INSPIRE as of 03 Mar 2017
  
%\cite{Vovchenko:2013viu}
\bibitem{Vovchenko:2013viu} 
  V.~Vovchenko, D.~Anchishkin and L.~P.~Csernai,
  %``Longitudinal fluctuations of the center of mass of the participants in heavy-ion collisions,''
  Phys.\ Rev.\ C {\bf 88}, no. 1, 014901 (2013)
%  doi:10.1103/PhysRevC.88.014901
  [arXiv:1306.5208 [nucl-th]].
  %%CITATION = doi:10.1103/PhysRevC.88.014901;%%
  %21 citations counted in INSPIRE as of 03 Mar 2017
  
%\cite{Olszewski:2015xba}
\bibitem{Olszewski:2015xba} 
  A.~Olszewski and W.~Broniowski,
  %``Multibin correlations in a superposition approach to relativistic heavy-ion collisions,''
  Phys.\ Rev.\ C {\bf 92}, no. 2, 024913 (2015)
%  doi:10.1103/PhysRevC.92.024913
  [arXiv:1502.05215 [nucl-th]].
  %%CITATION = doi:10.1103/PhysRevC.92.024913;%%
  %6 citations counted in INSPIRE as of 27 Feb 2017
  
%\cite{Casalderrey-Solana:2013sxa}
\bibitem{Casalderrey-Solana:2013sxa} 
  J.~Casalderrey-Solana, M.~P.~Heller, D.~Mateos and W.~van der Schee,
  %``Longitudinal Coherence in a Holographic Model of Asymmetric Collisions,''
  Phys.\ Rev.\ Lett.\  {\bf 112}, no. 22, 221602 (2014)
%  doi:10.1103/PhysRevLett.112.221602
  [arXiv:1312.2956 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevLett.112.221602;%%
  %30 citations counted in INSPIRE as of 03 Mar 2017
  
%\cite{Vechernin:2015upa}
\bibitem{Vechernin:2015upa} 
  V.~Vechernin,
  %``Forward–backward correlations between multiplicities in windows separated in azimuth and rapidity,''
  Nucl.\ Phys.\ A {\bf 939}, 21 (2015).
%  doi:10.1016/j.nuclphysa.2015.03.009
  %%CITATION = doi:10.1016/j.nuclphysa.2015.03.009;%%
  %11 citations counted in INSPIRE as of 27 Feb 2017
  
%\cite{Pang:2015zrq}
\bibitem{Pang:2015zrq} 
  L.~G.~Pang, H.~Petersen, G.~Y.~Qin, V.~Roy and X.~N.~Wang,
  %``Decorrelation of anisotropic flow along the longitudinal direction,''
  Eur.\ Phys.\ J.\ A {\bf 52}, no. 4, 97 (2016)
%  doi:10.1140/epja/i2016-16097-x
  [arXiv:1511.04131 [nucl-th]].
  %%CITATION = doi:10.1140/epja/i2016-16097-x;%%
  %19 citations counted in INSPIRE as of 03 Mar 2017
  
%\cite{Bialas:1976ed}
\bibitem{Bialas:1976ed} 
  A.~Bialas, M.~Bleszynski and W.~Czyz,
  %``Multiplicity Distributions in Nucleus-Nucleus Collisions at High-Energies,''
  Nucl.\ Phys.\ B {\bf 111}, 461 (1976).
%  doi:10.1016/0550-3213(76)90329-1
  %%CITATION = doi:10.1016/0550-3213(76)90329-1;%%
  %615 citations counted in INSPIRE as of 27 Feb 2017
  
%\cite{Bialas:2004su}
\bibitem{Bialas:2004su} 
  A.~Bialas and W.~Czyz,
  %``Wounded nucleon model and Deuteron-Gold collisions at RHIC,''
  Acta Phys.\ Polon.\ B {\bf 36}, 905 (2005)
  [hep-ph/0410265].
  %%CITATION = HEP-PH/0410265;%%
  %85 citations counted in INSPIRE as of 27 Feb 2017
  
%\cite{Gavin:2006xd}
\bibitem{Gavin:2006xd} 
  S.~Gavin and M.~Abdel-Aziz,
  %``Measuring Shear Viscosity Using Transverse Momentum Correlations in Relativistic Nuclear Collisions,''
  Phys.\ Rev.\ Lett.\  {\bf 97}, 162302 (2006)
%  doi:10.1103/PhysRevLett.97.162302
  [nucl-th/0606061].
  %%CITATION = doi:10.1103/PhysRevLett.97.162302;%%
  %123 citations counted in INSPIRE as of 27 Feb 2017
  
%\cite{Gavin:2016hmv}
\bibitem{Gavin:2016hmv} 
  S.~Gavin, G.~Moschelli and C.~Zin,
  %``Rapidity Correlation Structure in Nuclear Collisions,''
  Phys.\ Rev.\ C {\bf 94}, no. 2, 024921 (2016)
%  doi:10.1103/PhysRevC.94.024921
  [arXiv:1606.02692 [nucl-th]].
  %%CITATION = doi:10.1103/PhysRevC.94.024921;%%
  %6 citations counted in INSPIRE as of 27 Feb 2017
  

%\cite{Braun:2003fn}
\bibitem{Braun:2003fn} 
  M.~A.~Braun, R.~S.~Kolevatov, C.~Pajares and V.~V.~Vechernin,
  %``Correlations between multiplicities and average transverse momentum in the percolating color strings approach,''
  Eur.\ Phys.\ J.\ C {\bf 32}, 535 (2004)
%  doi:10.1140/epjc/s2003-01443-6
  [hep-ph/0307056].
  %%CITATION = doi:10.1140/epjc/s2003-01443-6;%%
  %57 citations counted in INSPIRE as of 17 Jun 2017  

  
%\cite{Bozek:2013sda}
\bibitem{Bozek:2013sda} 
  P.~Bozek, A.~Bzdak and V.~Skokov,
  %``The rapidity dependence of the average transverse momentum in p+Pb collisions at the LHC: the Color Glass Condensate versus hydrodynamics,''
  Phys.\ Lett.\ B {\bf 728}, 662 (2014)
%  doi:10.1016/j.physletb.2013.12.034
  [arXiv:1309.7358 [hep-ph]].
  %%CITATION = doi:10.1016/j.physletb.2013.12.034;%%
  %24 citations counted in INSPIRE as of 03 Mar 2017
  
%\cite{Gelis:2010nm}
\bibitem{Gelis:2010nm} 
  F.~Gelis, E.~Iancu, J.~Jalilian-Marian and R.~Venugopalan,
  %``The Color Glass Condensate,''
  Ann.\ Rev.\ Nucl.\ Part.\ Sci.\  {\bf 60}, 463 (2010)
%  doi:10.1146/annurev.nucl.010909.083629
  [arXiv:1002.0333 [hep-ph]].
  %%CITATION = doi:10.1146/annurev.nucl.010909.083629;%%
  %546 citations counted in INSPIRE as of 03 Mar 2017

%\cite{Blaizot:2016qgz}
\bibitem{Blaizot:2016qgz} 
  J.~P.~Blaizot,
  %``High Gluon Densities in Heavy Ions Collisions,''
  Rept.\ Prog.\ Phys.\  {\bf 80}, no. 3, 032301 (2017)
%  doi:10.1088/1361-6633/aa5435
  [arXiv:1607.04448 [hep-ph]].
  %%CITATION = doi:10.1088/1361-6633/aa5435;%%
  %1 citations counted in INSPIRE as of 03 Mar 2017
  
%\cite{Marquet:2006xm}
\bibitem{Marquet:2006xm} 
  C.~Marquet, G.~Soyez and B.~W.~Xiao,
  %``On the probability distribution of the stochastic saturation scale in QCD,''
  Phys.\ Lett.\ B {\bf 639}, 635 (2006)
%  doi:10.1016/j.physletb.2006.07.022
  [hep-ph/0606233].
  %%CITATION = doi:10.1016/j.physletb.2006.07.022;%%
  %40 citations counted in INSPIRE as of 05 Mar 2017
  
%\cite{McLerran:2015lta}
\bibitem{McLerran:2015lta} 
  L.~McLerran and M.~Praszalowicz,
  %``Fluctuations and the rapidity dependence of charged particles spectra in fixed centrality bins in p$A$ collisions,''
  Annals Phys.\  {\bf 372}, 215 (2016)
%  doi:10.1016/j.aop.2016.05.010
  [arXiv:1507.05976 [hep-ph]].
  %%CITATION = doi:10.1016/j.aop.2016.05.010;%%
  %9 citations counted in INSPIRE as of 05 Mar 2017
  
%\cite{McLerran:2015qxa}
\bibitem{McLerran:2015qxa} 
  L.~McLerran and P.~Tribedy,
  %``Intrinsic Fluctuations of the Proton Saturation Momentum Scale in High Multiplicity p+p Collisions,''
  Nucl.\ Phys.\ A {\bf 945}, 216 (2016)
%  doi:10.1016/j.nuclphysa.2015.10.008
  [arXiv:1508.03292 [hep-ph]].
  %%CITATION = doi:10.1016/j.nuclphysa.2015.10.008;%%
  %15 citations counted in INSPIRE as of 05 Mar 2017
  
%\cite{Mantysaari:2016jaz}
\bibitem{Mantysaari:2016jaz} 
  H.~Mäntysaari and B.~Schenke,
  %``Revealing proton shape fluctuations with incoherent diffraction at high energy,''
  Phys.\ Rev.\ D {\bf 94}, no. 3, 034042 (2016)
%  doi:10.1103/PhysRevD.94.034042
  [arXiv:1607.01711 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.94.034042;%%
  %7 citations counted in INSPIRE as of 05 Mar 2017
  
%\cite{Praszalowicz:2015dta}
\bibitem{Praszalowicz:2015dta} 
  M.~Praszalowicz and A.~Francuz,
  %``Geometrical Scaling in Inelastic Inclusive Particle Production at the LHC,''
  Phys.\ Rev.\ D {\bf 92}, no. 7, 074036 (2015)
%  doi:10.1103/PhysRevD.92.074036
  [arXiv:1507.08186 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.92.074036;%%
  %7 citations counted in INSPIRE as of 03 Mar 2017
  
%\cite{Dumitru:2001ux}
\bibitem{Dumitru:2001ux} 
  A.~Dumitru and L.~D.~McLerran,
  %``How protons shatter colored glass,''
  Nucl.\ Phys.\ A {\bf 700}, 492 (2002)
%  doi:10.1016/S0375-9474(01)01301-X
  [hep-ph/0105268].
  %%CITATION = doi:10.1016/S0375-9474(01)01301-X;%%
  %148 citations counted in INSPIRE as of 03 Mar 2017
  
%\cite{Duraes:2015qoa}
\bibitem{Duraes:2015qoa} 
  F.~O.~Duraes, A.~V.~Giannini, V.~P.~Goncalves and F.~S.~Navarra,
  %``Rapidity dependence of the average transverse momentum in hadronic collisions,''
  Phys.\ Rev.\ C {\bf 94}, no. 2, 024917 (2016)
%  doi:10.1103/PhysRevC.94.024917
  [arXiv:1510.04737 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevC.94.024917;%%
  %3 citations counted in INSPIRE as of 03 Mar 2017
  
%\cite{Deja:2017dqh}
\bibitem{Deja:2017dqh} 
  K.~Deja and K.~Kutak,
  %``Rapidity dependence of the average transverse momentum in p+p and p+Pb collisions revisited,''
  Phys.\ Rev.\ D {\bf 95}, no. 11, 114027 (2017)
%  doi:10.1103/PhysRevD.95.114027
  [arXiv:1703.07959 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.95.114027;%%
  
%\cite{DiFrancesco:2016srj}
\bibitem{DiFrancesco:2016srj} 
  P.~Di Francesco, M.~Guilbaud, M.~Luzum and J.~Y.~Ollitrault,
  %``Systematic procedure for analyzing cumulants at any order,''
  Phys.\ Rev.\ C {\bf 95}, no. 4, 044911 (2017)
%  doi:10.1103/PhysRevC.95.044911
  [arXiv:1612.05634 [nucl-th]].
  %%CITATION = doi:10.1103/PhysRevC.95.044911;%%
  %7 citations counted in INSPIRE as of 10 Sep 2017

%\cite{Bhalerao:2014mua}
\bibitem{Bhalerao:2014mua} 
  R.~S.~Bhalerao, J.~Y.~Ollitrault, S.~Pal and D.~Teaney,
  %``Principal component analysis of event-by-event fluctuations,''
  Phys.\ Rev.\ Lett.\  {\bf 114}, no. 15, 152301 (2015)
%  doi:10.1103/PhysRevLett.114.152301
  [arXiv:1410.7739 [nucl-th]].
  %%CITATION = doi:10.1103/PhysRevLett.114.152301;%%
  %28 citations counted in INSPIRE as of 27 Feb 2017

\end{thebibliography}

\documentclass[10pt,notitlepage,nofootinbib,twocolumn]{revtex4-1}
\usepackage{graphicx,color}
\usepackage{hyperref}
\usepackage{amssymb}

\begin{document}

\title{Rapidity dependence of transverse-momentum multiplicity correlations}
\author{Adam Bzdak}
\email[E-Mail:]{bzdak@fis.agh.edu.pl}

\affiliation{AGH University of Science and Technology,\\
Faculty of Physics and Applied Computer Science,\\
30-059 Krak\'ow, Poland}

\begin{abstract}
Following previous work \cite{Bzdak:2012tp}, we propose to analyze the 
rapidity dependence of transverse momentum and transverse-momentum multiplicity correlations.
We demonstrate that the orthogonal polynomial expansion of the latter has the potential to discriminate between models of particle production.
\end{abstract}

\maketitle

\section{Introduction}
\label{sec: introduction}

One of the central problems in high-energy hadronic collisions
is to understand the longitudinal structure of systems created in
proton-proton (p+p), proton-nucleus (p+A) and nucleus-nucleus (A+A)
collisions.

Not long ago, it was argued that an event-by-event long-range fluctuation of
the fireball rapidity distribution results in rather peculiar two- and
multi-particle rapidity correlations \cite{Bzdak:2012tp,Bzdak:2015dja}.
Recent measurement by the ATLAS Collaboration at the LHC \cite%
{Aaboud:2016jnr} revealed new and rather unexpected scaling results on
asymmetric rapidity fluctuations in p+p, p+A and A+A interactions.
Recently, this problem has drawn a noticeable theoretical \cite%
{Jia:2015jga,Bozek:2015tca,Monnai:2015sca,Broniowski:2015oif,Schenke:2016ksl,Bzd-Dus,Ke:2016jrd,He:2017laa}
and experimental \cite{Aaboud:2016jnr,star:a1,alice:a1} interest, see also %
\cite%
{Bozek:2010vz,Bialas:2011xk,Bialas:2011bz,Csernai:2012mh,Pang:2012he,Vovchenko:2013viu,Olszewski:2015xba,Casalderrey-Solana:2013sxa,Vechernin:2015upa,Pang:2015zrq}
for recent related studies.

To summarize the main idea, the single-particle rapidity distribution in
each event, $N(y)$, can be written as
\begin{equation}
\frac{N(y)}{\left\langle N(y)\right\rangle }=1+a_{0}+a_{1}y+...,
\label{eq:N(y)}
\end{equation}%
where $a_{0}$ describes the rapidity independent fluctuation of the
fireball. $a_{1}$ represents the fluctuating long-range
forward-backward rapidity asymmetry.\footnote{%
We are not interested in statistical fluctuations, which can also generate
nonzero values of $a_i$. These are removed by measuring correlation
functions.} This coefficient can be driven for example by the difference in the
number of left- and right-going sources of particles, e.g., wounded nucleons 
\cite{Bialas:1976ed,Bialas:2004su}.  
$\left\langle N(y)\right\rangle $ is the average rapidity distribution in a given
centrality class. By definition $\left\langle
a_{i}\right\rangle =0$.

It is straightforward to calculate the two-particle rapidity correlation %
\cite{Bzdak:2012tp} 
\begin{equation}
\frac{C(y_{1},y_{2})}{\left\langle N(y_{1})\right\rangle \left\langle
N(y_{2})\right\rangle }=\left\langle a_{0}^{2}\right\rangle +\left\langle
a_{1}^{2}\right\rangle y_{1}y_{2}+...  \label{eq:C-NN}
\end{equation} 
where\footnote{For clarity we skip $\left\langle a_{0}a_{1}\right\rangle$, 
which vanishes in symmetric (e.g. p+p) collisions.}
\begin{equation}
C(y_{1},y_{2})=\left\langle N(y_{1})N(y_{2})\right\rangle -\left\langle
N(y_{1})\right\rangle \left\langle N(y_{2})\right\rangle .
\end{equation}
As seen in Eq. (\ref{eq:C-NN}), the long-range fluctuation of the fireball rapidity distribution, 
parameterized by fluctuating $a_i$, results in rather nontrivial correlations. The first term
corresponds to a well-known rapidity independent multiplicity fluctuation,
and it can be driven by, e.g., the impact parameter or volume
fluctuation. The second term, $\sim y_{1}y_{2}$, is related to the
fluctuating forward-backward asymmetry in rapidity. In the
wounded nucleon model \cite{Bialas:1976ed,Bialas:2004su} $\langle
a_{1}^{2}\rangle \sim \langle (w_{L}-w_{R})^{2}\rangle ,$
where $w_{L(R)}$ is the number of left(right)-going wounded nucleons \cite{Bzdak:2012tp}.
Recently, the ATLAS Collaboration observed $\langle
a_{1}^{2}\rangle y_{1}y_{2}$ in the two-particle rapidity correlation
functions measured in p+p, p+Pb and Pb+Pb collisions \cite{Aaboud:2016jnr}.
They found that at a given event multiplicity $N_{\mathrm{ch}}$, $%
\langle a_{1}^{2}\rangle $ approximately scales with $1/N_{%
\mathrm{ch}}$ and numerically is very similar for all colliding systems.%
\footnote{%
One would expect rather different results in, e.g., peripheral Pb+Pb
and ultracentral p+p collisions ($N_{\mathrm{ch}} \sim 150$
\cite{Aaboud:2016jnr}). The ATLAS result suggests that the number of
particle sources (at a given $N_{\mathrm{ch}}$) and their fluctuations are
actually similar in all measured systems.} This surprising result still
calls for a quantitative explanation.

\section{Transverse-momentum multiplicity correlations}

It is proposed here to analyze, in a similar way, the rapidity dependence of
transverse momentum and especially transverse-momentum multiplicity correlations. 

Analogously to Eq. (\ref{eq:N(y)}) we have 
\begin{equation}
\frac{P_{t}(y)}{\left\langle P_{t}(y)\right\rangle }=1+b_{0}+b_{1}y+...,
\end{equation}%
where $P_{t}(y)$ is the average (in one event) transverse momentum of
particles in a given rapidity bin $y$
\begin{equation}
P_{t}(y)=\frac{1}{N}\sum\nolimits_{i=1}^{N}p_{t}^{(i)},
\end{equation}
where $N$ is the number of particles (in a given event) at~$y$. Here $%
p_{t}^{(i)}$ is the transverse momentum magnitude of the $i$-th particle. $%
\left\langle P_{t}(y)\right\rangle $ is the average of $P_{t}(y)$ over many
events in a given centrality class. We note that $\left\langle b_{i}\right\rangle =0$, 
in close analogy to the $a_{i}$ coefficients.

The transverse momentum correlation function (studied extensively in the
literature for rather different reasons, see, e.g., \cite%
{Gavin:2006xd,Gavin:2016hmv,Braun:2003fn}) reads 
\begin{equation}
\frac{C_{[P,P]}(y_{1},y_{2})}{\left\langle P_{t}(y_{1})\right\rangle
\left\langle P_{t}(y_{2})\right\rangle }=\left\langle b_{0}^{2}\right\rangle
+\left\langle b_{1}^{2}\right\rangle y_{1}y_{2}+...,  \label{eq:C-PP}
\end{equation}
where 
\begin{equation}
C_{[P,P]}(y_{1},y_{2})\equiv \left\langle
P_{t}(y_{1})P_{t}(y_{2})\right\rangle -\left\langle
P_{t}(y_{1})\right\rangle \left\langle P_{t}(y_{2})\right\rangle.
\end{equation}

The first term in Eq. (\ref{eq:C-PP}) describes an event-by-event rapidity
independent transverse momentum fluctuation. This could be driven for
example by an event-by-event long-range multiplicity fluctuation (if event
multiplicity is correlated with $P_{t}$). The second term describes the
forward-backward rapidity asymmetric transverse momentum fluctuation. 
A possible source of this effect is the forward-backward fireball multiplicity
fluctuation.

It would be especially interesting to measure an event-by-event relation
between $a_{i}$ and $b_{i}$ coefficients. In order to do this, one can
construct a simple correlation function 
\begin{equation}
C_{[N,P]}(y_{1},y_{2})\equiv \left\langle N(y_{1})P_{t}(y_{2})\right\rangle
-\left\langle N(y_{1})\right\rangle \left\langle P_{t}(y_{2})\right\rangle ,
\end{equation}
witch correlates multiplicity and transverse momentum, see, e.g., \cite{Braun:2003fn}. This results in 
\begin{equation}
\frac{C_{[N,P]}(y_{1},y_{2})}{\left\langle N(y_{1})\right\rangle
\left\langle P_{t}(y_{2})\right\rangle }=\left\langle
a_{0}b_{0}\right\rangle +\left\langle a_{1}b_{1}\right\rangle y_{1}y_{2}+...
\end{equation}

The meaning of mixed coefficients $\left\langle a_{i}b_{k}\right\rangle $ is
easy to understand.  
The first term describes the relation between rapidity independent
fluctuation of multiplicity and transverse momentum. The second term is
particularly interesting and describes how rapidity asymmetry in
multiplicity is related to rapidity asymmetry of transverse momentum. If the
particle multiplicity and $P_{t}$ are not correlated then $\left\langle
a_{i}b_{k}\right\rangle =\left\langle a_{i}\right\rangle \left\langle
b_{k}\right\rangle =0$.

In general, the above correlation functions can be expanded in terms of the
orthogonal polynomials \cite{Bzdak:2012tp}. For example 
\begin{equation}
\frac{C_{[N,P]}(y_{1},y_{2})}{\left\langle N(y_{1})\right\rangle
\left\langle P_{t}(y_{2})\right\rangle }=\sum\nolimits_{i,k}\left\langle
a_{i}b_{i}\right\rangle T_{i}(y_{1})T_{k}(y_{2}),
\end{equation}%
with $T_{i}$ being, e.g., the Chebyshev or the Legendre polynomials \cite%
{Bzdak:2012tp,Jia:2015jga}, and analogously for Eqs. (\ref{eq:C-NN}) and (\ref{eq:C-PP}).

\section{Discussion and conclusions}

Several comments are in order.

Consider a set of events with $a_{1}>0$, i.e., the fireball multiplicity is larger
for positive $y$, $N(y) \sim a_{1}y$. The question is what is the rapidity
dependence of the transverse momentum in this case. If $P_{t}$ is also
larger for positive $y$ then $b_{1}>0$ and thus $\langle
a_{1}b_{1}\rangle >0$. This scenario is expected in a typical
hydrodynamical framework, see, e.g., \cite{Bozek:2013sda}.

For example, in the color glass condensate (CGC) framework \cite%
{Gelis:2010nm,Blaizot:2016qgz} one could expect a rather different
conclusion. Consider a proton-proton event, where the two protons are
characterized by different saturation scales, $Q_{1}$ and $Q_{2}$. The
importance of such fluctuations was recently discussed in Refs. \cite%
{Marquet:2006xm,McLerran:2015lta,McLerran:2015qxa,Mantysaari:2016jaz,Bzd-Dus}%
. Here $Q_{1}^{2}=Q_{0,1}^{2}e^{+\lambda y}$ and $Q_{2}^{2}=Q_{0,2}^{2}e^{-%
\lambda y}$ with $\lambda \sim 0.3$, see, e.g., \cite{Praszalowicz:2015dta}.
We choose $Q_{0,1}>Q_{0,2}$ so that in a given rapidity bin, say $|y|<2$, $%
Q_{1}>Q_{2}$, resulting in rapidity asymmetric $N(y)$. 
In this case \cite{Dumitru:2001ux,Bozek:2013sda} 
\begin{eqnarray}
N(y) &\sim &S_{t}Q_{2}^{2}\left[ 2+\ln \left( Q_{1}^{2}/Q_{2}^{2}\right) %
\right] , \\[5pt]
P_{t}(y) &\sim &\frac{2Q_{1}-\frac{2}{3}Q_{2}}{1+\ln \left(
Q_{1}/Q_{2}\right) },
\end{eqnarray}
that is, in CGC the multiplicity is driven by the smaller scale in
contrast to the transverse momentum controlled by the larger one \cite{Dumitru:2001ux}. 
Since $Q_{1}^{2}\sim e^{+\lambda y}$ and 
$Q_{2}^{2}\sim e^{-\lambda y}$, the multiplicity and the transverse momentum
rapidity asymmetries have different signs. If $N(y)$ is growing with
rapidity, then $P_{t}(y)$ is decreasing with $y$. Consequently $a_{1} \gtrless 0$
means $b_{1} \lessgtr 0$ and $\left\langle a_{1}b_{1}\right\rangle <0$. Clearly, this
observation should be treated with caution and more detailed calculations are
warranted, see, e.g., \cite{Duraes:2015qoa,Deja:2017dqh}. The sole purpose of this
exercise was to demonstrate that the sign of $\left\langle
a_{1}b_{1}\right\rangle $ is not at all obvious, and could potentially discriminate 
between different models of particle production. 

As discussed earlier, the ATLAS Collaboration reported a surprising
scaling of $\langle a_{1}^{2}\rangle $ in p+p, p+Pb and Pb+Pb
collisions \cite{Aaboud:2016jnr}. At a given event multiplicity 
$N_{\mathrm{ch}}$, $\langle a_{1}^{2}\rangle $ scales with $1/N_{\rm ch}$ and is
quantitatively very similar for all three systems. It would be
very interesting to see if $\langle b_{1}^{2}\rangle $ and $%
\left\langle a_{1}b_{1}\right\rangle $ satisfy similar scaling.

Obviously, it would be also desired to study higher order
correlation functions \cite{Bzdak:2015dja,DiFrancesco:2016srj}.

An alternative way to analyze the above correlation functions is 
the principal component analysis, discussed in Ref. \cite{Bhalerao:2014mua}.

In conclusion, it is proposed to analyze the rapidity dependence of
transverse momentum and in particular transverse-momentum multiplicity
correlation functions using the orthogonal polynomial
expansion. A careful study of the coefficients $\langle
a_{i}^{2}\rangle ,$ $\langle b_{i}^{2}\rangle $ and $\langle a_{i}b_{k}\rangle $ 
could potentially discriminate between different 
models of particle production, and reveal detailed information on the longitudinal 
structure of systems created in p+p, p+A and A+A collisions.

\bigskip

\vspace{\baselineskip} 
\noindent\textbf{Acknowledgments} 
\newline
{} 
We thank Piotr Bo\.zek and Volker Koch for useful comments. 
This work is supported by the
Ministry of Science and Higher Education (MNiSW) and by the National Science
Centre, Grant No. DEC-2014/15/B/ST2/00175, and in part by
DEC-2013/09/B/ST2/00497.

\end{document}